OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (anonymous):

@madhu.mukherjee.946 @hartnn

OpenStudy (midhun.madhu1987):

Perimeter of Rectangle P = 2 * ( l + b ) Let length, l = 7x and Breadth, b = 5x P = 2 * ( l + b ) 96 = 2 * ( 7x + 5x ) Solve for x. Then Find length and breadth.

OpenStudy (anonymous):

i'm getting C is that right?

OpenStudy (midhun.madhu1987):

Can you let me know how you got that answer?

OpenStudy (anonymous):

7x+5x=12x*2=24

OpenStudy (midhun.madhu1987):

96 = 2 * ( 7x + 5x ) 12x * 2 = 96 24x = 96 What is the value of x?

OpenStudy (anonymous):

do you divide?

OpenStudy (anonymous):

no you dont ok

OpenStudy (anonymous):

i'm still confused on what to do next?

OpenStudy (midhun.madhu1987):

24x = 96 Divide both sides by 24. What you get?

OpenStudy (anonymous):

1 and 0.25

OpenStudy (midhun.madhu1987):

24x = 96 so, x = 96 / 24 = 4 we got the value of x = 4. Now find length = 7x = 7 * 4 = 28 Breadth = 5x = 5* 4 = 20

OpenStudy (anonymous):

ohhh ok it was B all along

OpenStudy (midhun.madhu1987):

yep :)